Solve using substitution or
Elimination.
Y=2x-1 and 2x + 2y = 22

Answer by Alan3354(69443)

Solve using substitution or
Elimination.
Y=2x-1 and 2x + 2y = 22
==========================
y=2x-1
x + y = 11
---
x + 2x-1 = 11
3x = 12
x = 4